OpenStudy (anonymous):

Solve, using factoring: 8x2 = 16x {0} {-8, -16} {0, 8} {0, 2}

hero (hero):

8x^2-16x = 0 8x(x-2) = 0 x = 0,2

OpenStudy (anonymous):

8x2=16x 8x2-16x=16x-16x 8x2-16x=0 8x(x-2)=0 x=0,2

OpenStudy (anonymous):

0 and 2